Metallic zinc reacts with aqueous HCl.

Zn(s) + 2 HCl(aq) → ZnCl2(aq) + H2(g)

What volume of 2.20 M HCl, in milliliters, is required to convert 12.8 g of Zn completely to products?


mL HCl
1
Expert's answer
2018-10-11T08:17:09-0400

12.8 g x

Zn + 2 HCl → ZnCl2 + H2

65 2 moles

X = 12.8*2/65 = 0.39 moles

CM = n/V

V = n/ CM = 0.39/2.20 = 0.179 L

V = 179 ml